OpenStudy (jennychan12):

indefinite integral of x^3e^x^2dx ?

OpenStudy (jennychan12):

\[\int\limits x^3e^{x^2}dx\] ? I know it's integration by parts...

OpenStudy (jennychan12):

would you do u = e^x^2 and dv = x^3 ?

OpenStudy (anonymous):

try letting u=x^3 and dv = e^(x^2)

OpenStudy (anonymous):

sorry other way around lol dv = x^3 and u = e^(x^2)

OpenStudy (jennychan12):

i did that before... so would you have to do integration by parts like two or three times?

OpenStudy (anonymous):

so v = 1/4 x^4 and du = 2xe^(x^2) dx

OpenStudy (anonymous):

then using the equation uv - \[\int\limits_{}^{} \] vdu you can use a U-substitution for that integral

OpenStudy (anonymous):

lol bad formatting its uv - integral of vdu

OpenStudy (jennychan12):

okay, so putting all that in, i get \[\int\limits x^3 e^{x^2} dx = \frac{ e^{x^2}x^4 }{ 4 } - 2 \int\limits x^4e^{x^2}dx\]

OpenStudy (jennychan12):

how would you go from there?

OpenStudy (anonymous):

sorry mistake do a u-sub first so let u =x^2

OpenStudy (anonymous):

i dont know if this is correct but. integral (x^3)(e^x^2)dx u = x^2 du = 2xdx du/2x = dx integral x(x^2)(e^u)(du/2x) so you would get intl u(e^u)du so...that would proll make it a little better. for IBP

OpenStudy (anonymous):

once you do the u-sub it will become \[\int\limits_{}^{} ue^u du \]

OpenStudy (anonymous):

and just use integration by parts from theere

OpenStudy (jennychan12):

i did the u = x^2 and i got \[ 2 \int\limits x^4e^{x^2}dx = 2(\frac{ x^7 }{ 3 } - \frac{ 4x^5 }{ 15 }) +C\]

OpenStudy (anonymous):

so u = x^2, du=2xdx so xdx=du/2 \[\int\limits_{}^{} x^2 e^u du\] and since u =x^2 you replace that x^2 with a u

OpenStudy (anonymous):

my bad theres also a 1/2 in that integral

OpenStudy (anonymous):

so it just becomes \[\frac{ 1 }{ 2 } \int\limits_{}^{}ue^u du\] and you use integration by parts from here

OpenStudy (raden):

int (1/2 u e^u) du use the tabular integration by part